I need help in finding the slope for (0,-9) and (-4,0) thanks
Found 2 solutions by checkley77, mathbath:
Answer by checkley77(12844)

(0,-9) and (-4,0)
Slope=(y2-y1)/(x2-x1)
Slope=(0+9)/(-4-0)
Slope=9/-4=-9/4=-2.25 ans.

Answer by mathbath(13)

The two co-ordinates in question are (0,-9) and (-4,0).
Now these two co-ordinates exist on the same straight line.
We know in general if a line is perfectly horizontal there is no slope to it.
If the line is perfectly vertical the slope of the line is infinity.
If the right end of the line is lower than the left end the slope is negative. This means the line is dropping
If the right end of the line is higher than the left end the slope is positive. This means the line is rising.
If (x1,y1) and (x2,y2) are any two co-ordinates on the line the slope "m" of the line is calculated by,
m=(y2-y1)/(x2-x1)
Substituting the values from the current problem,
x1=0, y1=-9
x2=-4, y2=0
m = ( 0 - (-9) )/( (-4) - 0 )
m = ( 0 + 9 )/( -4 )
m = 9/(-4)
m = -9/4
So the slope of the line is -9/4.
Since the slope has a negative value the line is dropping.